OpenStudy (anonymous):

Help!! will give medals :) Given:∠1≅∠2 Prove:m∠3+m∠4=90 Statements Reason 1)∠1≅∠2 Given 2)??????? ??????? 3)????? ?????? 4)m∠3+m∠4=90 ?????

OpenStudy (anonymous):



OpenStudy (anonymous):

i think 2 is l is perpendicular to m 3 is angle 3 is congruent to angle 4 by theorem 3.10 4 is definition of complementarty angles

OpenStudy (anonymous):

@uri @kelliegirl33 @skullpatrol

OpenStudy (anonymous):

@Loser66 @GoldPhenoix

OpenStudy (primeralph):

So you ask questions too.

OpenStudy (primeralph):

To prove.

OpenStudy (anonymous):

Based on what's given, it's impossible to prove it

OpenStudy (primeralph):

@FutureMathProfessor Actually, it is. Look at the picture.

OpenStudy (anonymous):

No it's not

OpenStudy (primeralph):

If a1 = a2, then a1 + a2 = 180 = 2(a1) = 180. a1 = a2 = 90 Then a5 can only be 90 because the line is a perpend, leaving the rest to be 360-(90+180). So a3 and a4 share 90.

OpenStudy (anonymous):

One cannot justify If a1 = a2, then a1 + a2 = 180

OpenStudy (primeralph):

@FutureMathProfessor Ever heard of "sum of angles on a straight line = 180?"

OpenStudy (anonymous):

One does not know that the line is straight

OpenStudy (anonymous):

@primeralph could you help me explain it in that format

OpenStudy (anonymous):

Very poorly written problem in my opinion

OpenStudy (primeralph):

@FutureMathProfessor Go home, you're drunk.